OpenStudy (anonymous):

$\left\{\left|x\right|<2:\sqrt{\cos x}\cos \left(300x\right)-.1+\sqrt{\left|x\right|}\right\}$

OpenStudy (unklerhaukus):

\[\left\{\left|x\right|<2:\sqrt{\cos x}\cos \left(300x\right)-.1+\sqrt{\left|x\right|}\right\}\]

OpenStudy (anonymous):

Ummm. Whats the question?